OpenStudy (anonymous):

Help!!!!!! Solve for x : 2log(base3) x-2 + 2log(base3) x-4 = 0

OpenStudy (anonymous):

x-2=(1/(x-4)) solve to get x

OpenStudy (anonymous):

\[2\log_{3} x-2 + 2\log_{3}x-4 = 0\]

OpenStudy (anonymous):

answer please!!!!!!!!

OpenStudy (shubhamsrg):

no buddy,,.you solve please..i mean try yourself..we may correct you whenever required..

OpenStudy (anonymous):

is (x-2)(x-4)=1 the answer?

OpenStudy (anonymous):

Ans. 3+root2 , 3

OpenStudy (anonymous):

x^2-6x+7=0

OpenStudy (anonymous):

hey nitz have you posted only x^2 - 6x +7 = 0

OpenStudy (anonymous):

because i am not able to see any other thing here.